The problem asks me to solve usling any method.
5x+6y=14
y=4x-17
I am not sure which method to use. I tried both but did not come out with the same answer both times! Thanks for your help!

Answer by rapaljer(4671)

Because one of the equations is in the form of y=___ (or also x=____), I think the substitution method is the easiest way.

Rewrite the first equation
5x + 6y = 14, and leave a blank space in place of the y.
5x + 6(_____) = 14.

In that blank space, substitute the value y from the second equation, that is, 4x-17.

5x + 6(_____) = 14
5x + 6(4x-17) = 14 Now just solve for x.
5x + 24x - 102 = 14
29x - 102 = 14
29x - 102 + 102 = 14 + 102
29x = 116 Divide both sides by 29
x= 4.

Now, the easiest place to find y is in the y=___ equation.
y=4x- 17
y=4(4) - 17
y = -1
Check in the OTHER equation:
5x+ 6y = 14
5(4) + 6(-1) = 14
20 - 6 = 14
It checks!!

R^2 at SCC